OpenStudy (anonymous):

9p-4p+6=7p-3p step by step plzz

OpenStudy (saifoo.khan):

9p-4p+6=7p-3p 5p + 6 = 4p 5p - 4p = - 6 p = -6

OpenStudy (anonymous):

combine like terms, then get p isolated.

OpenStudy (anonymous):

9p-4p+6 = 7p-3p 9p-4p-7p+3p = -6 p = -6

OpenStudy (agreene):

(9-4)p+6 = (7-3)p 5p+6=4p subtract 4p and 6 from both sides (5-4)p = -6 p=-6

OpenStudy (anonymous):

Give them a chance to do it on their owwwwwwwn